# EHRAgent: Code Empowers Large Language Models for Complex Tabular Reasoning on Electronic Health Records

### v1

@@ 3.3 Error Tracing via Rubber Duck Debugging
The interaction ends either when a ‘TERMINATE’ signal appears in the generated messages or when $t$ reaches a pre-defined threshold $T$.
3.4 Plan Refinement with Long-term Memory
where $\arg\operatorname{TopK}{\max}(\cdot)$ identifies the indices of the top $K$ elements with the highest values from $\mathcal{L}$, and $\operatorname{sim}(\cdot,\cdot)$ calculates the similarity between two questions, employing negative Levenshtein distance as the similarity metric. Subsequent to this retrieval process, the newly acquired $K$-shot examples $\mathcal{E}(q)$ replace the originally predefined examples $E_{1},\cdots,E_{K}$ in Eq. (1). This updated set of examples serves to reformulate the prompt, guiding the LLM agent in plan refinement:
@@ 4.1 Experiment Setup
Tasks and Datasets. We evaluate EHRAgent on two publicly available structured EHR datasets, MIMIC-III (Johnson et al., 2016) and eICU (Pollard et al., 2018), from a question-answering benchmark (Lee et al., 2022) on EHRs. These questions originate from real-world clinical needs and cover a wide range of tabular queries commonly posed within EHRs. During data pre-processing stage, we create EHR question-answering pairs by considering text queries as questions and executing SQL commands in the database to automatically generate the corresponding ground-truth answers. Throughout this process, we filter out samples containing unexecutable SQL commands or yielding empty results. In total, the final dataset involves 13.5 tables and 580 examples per dataset, with an average of 2.13 tables required to answer each question. Dataset statistics are available in Table 1. More details can be found in Appendix A.
Table 1: Dataset statistics of MIMIC-III and eICU.
Dataset # Examples # Table # Row/Table # Table/Q
MIMIC-III 580 17 81k 2.52
eICU 580 10 152k 1.74
Total 1160 27 108k 2.13
@@
Tool Sets. To enable LLMs in complex operations such as calculations and information retrieval, we integrate external tools in EHRAgent during the interaction with EHRs. Our toolkit can be easily expanded with natural language tool function definitions in a plug-and-play manner. Toolset details are available in Appendix B.1.
Table 2: Main results of success rate (i.e., SR.) and completion rate (i.e., CR.) on MIMIC-III and eICU datasets. The complexity of questions increases from Level I (the simplest) to Level IV (the most difficult).
Dataset MIMIC-III eICU
Complexity Level I II III IV All I II III All
Metrics SR. SR. CR. SR. SR. CR.
CoT (Wei et al., 2022) 29.33 12.88 3.08 2.11 9.58 38.23 26.73 33.00 8.33 27.34 65.65
Chameleon (Lu et al., 2023) 38.67 14.11 4.62 4.21 12.77 42.76 31.09 34.68 16.67 35.06 83.41
ReAct (Yao et al., 2023b) 34.67 12.27 3.85 2.11 10.38 25.92 27.82 34.24 15.38 33.33 73.68
LLM2SQL (Nan et al., 2023) 23.68 10.64 6.98 4.83 13.10 44.83 20.48 25.13 12.50 23.28 51.72
AutoGen (Wu et al., 2023b) 36.00 28.13 15.33 11.11 22.49 61.47 42.77 40.70 18.75 40.69 86.21
EHRAgent (Ours) 71.58 66.34 49.70 49.14 58.97 85.86 54.82 53.52 25.00 53.10 91.72
Baselines. We compare EHRAgent with the following LLMs-based planning, tool using, and coding baselines. We summarize and compare their key designs with EHRAgent in Table 4 in Appendix B.2.
$\diamond$ CoT (Wei et al., 2022): It enhances the complex reasoning capabilities of original LLMs by generating a series of intermediate reasoning steps.
$\diamond$ Chameleon (Lu et al., 2023): It employs LLMs as controllers and integrates a set of plug-and-play modules, enabling enhanced reasoning and problem-solving across diverse tasks.
$\diamond$ ReAct (Yao et al., 2023b): It integrates reasoning with tool-use by guiding LLMs to generate intermediate verbal reasoning traces and tool commands.
$\diamond$ LLM2SQL (Nan et al., 2023): It augments LLMs with a code interface to generate SQL queries for retrieving information from EHRs for question answering.
$\diamond$ AutoGen (Wu et al., 2023b): It unifies LLM-based agent workflows as multi-agent conversations and uses the code interface to encode interactions between agents and environments.
@@
Evaluation Protocol. Our primary evaluation metric is the success rate, quantifying the percentage of queries that the model successfully handles. Furthermore, we assess the completion rate, which represents the percentage of queries that the model is able to generate executable plans (even not yield correct results). We categorize input queries into various complexity levels (I-IV) based on the number of tables involved in solution generation (see details in Appendix A.2).
@@
Implementation Details. We employ GPT-4 (OpenAI, 2023) as the base LLM model for all experiments. We set the temperature parameter ($t$) to 0 when making API calls to GPT-4 to eliminate randomness and set the pre-defined threshold ($T$) to 10. Due to the maximum length limitations of input context in baselines (e.g., ReAct and Chameleon), we use the same initial four-shot demonstrations ($K=4$) for all baselines and EHRAgent to ensure a fair comparison. Additional implementation details with prompt templates are available in Appendix B.
@@ 4.2 Main Results
Table 2 summarizes the experimental results of EHRAgent and baselines on the two datasets. From the results, we have the following observations:
(1) EHRAgent significantly outperforms all the baselines on both datasets with a performance gain of 36.48% and 12.41%, respectively. This indicates the efficacy of our key designs, namely interactive coding with environment feedback and domain knowledge injection, as they gradually refine the generated code and provide sufficient background knowledge during the planning process.
(2) CoT and Chameleon both neglect environment feedback and cannot adaptively refine their planning processes. Such deficiencies hinder their performance in EHR question-answering scenarios, as the success rate for these methods on two datasets are all below 40%.
(3) ReAct considers environment feedback but is restricted to tool-generated error messages. Consequently, it potentially overlooks the overall planning process. Moreover, it lacks a code interface, which prevents it from efficient action planning, and results in lengthy context execution and lower completion rates.
(4) LLM2SQL leverages LLM to directly generate SQL queries for EHR question-answering tasks. However, the gain is rather limited, as the LLM still struggles to generate high-quality SQL codes for execution. Besides, the absence of a dedicated code debugging module further impedes its overall performance for this challenging task.
(5) AutoGen presents a notable performance gain over other baselines, as it leverages code interfaces and considers the errors from the coding environment, leading to a large improvement in the completion rate. However, as it does not model medical knowledge or identify inherent error patterns in the current code, the performance is still sub-optimal.
@@ 4.3 Quantitative Analysis
Ablation Studies. Our ablation studies on MIMIC-III (Table 3) demonstrate the effectiveness of all four components in EHRAgent. Interactive coding is the most significant contributor across all complexity levels, which highlights the importance of code generation in planning and environmental interaction for refinement. In addition, more challenging tasks benefits more from knowledge integration, indicating that comprehensive understanding of EHRs facilitates the complex multi-tabular reasoning in effective schema linking and reference (e.g., tables, columns, and condition values) identification. Detailed analysis with additional results on eICU is available in Appendix C.1.
Complexity level I II III IV All
Metrics SR. SR. CR.
EHRAgent 71.58 66.34 49.70 49.14 58.97 85.86
 w/o medical knowledge 68.42 33.33 29.63 20.00 33.66 69.22
 w/o interactive coding 45.33 23.90 20.97 13.33 24.55 62.14
 w/o debugging 55.00 38.46 41.67 35.71 42.86 77.19
 w/o long-term memory 65.96 54.46 37.13 42.74 51.73 83.42
Effect of Question Complexity. We take a closer look at the model performance by considering multi-dimensional measurements of question complexity, exhibited in Figure 3. Although the performances of both EHRAgent and the baselines generally decrease with an increase in task complexity (either quantified as more elements in queries or more columns in solutions), EHRAgent consistently outperforms all the baselines at various levels of difficulty. Appendix D.1 includes additional analysis on the effect of various question complexities.
@@
Sample Efficiency. Figure 4 illustrates the model performance w.r.t.  number of demonstrations for EHRAgent and the strongest baseline, AutoGen. Comparing to supervised learning (e.g., text-to-SQL (Wang et al., 2020; Raghavan et al., 2021; Lee et al., 2022)) that requires extensive training on over 10K samples with detailed annotations (e.g., SQL code), LLM agents enable complex tabular reasoning using a few demonstrations only. One interesting finding is that as the number of examples increases, both the success and completion rate of AutoGen tend to decrease, mainly due to the context limitation of LLMs. On the contrary, the performance of EHRAgent remains stable with more demonstrations, which may benefit from its integration of a ‘rubber duck’ debugging module and the adaptive mechanism for selecting the most relevant demonstrations.
@@ 4.4 Case Study
Figure 5 showcases a comparative case study of the interactive coding process between AutoGen and EHRAgent for the same given query. When executed with error feedback, AutoGen directly sends back the original error messages, making slight modifications (e.g., changing the surface string of the arguments) without reasoning the root cause of the error. In contrast, EHRAgent can identify the underlying causes of the errors through interactive coding and debugging processes. It successfully discovers the underlying error causes (taking into account case-sensitivity), facilitating accurate code refinement. Extensive case studies showcasing the effectiveness of different components in EHRAgent are available in Appendix E.
4.5 Error Analysis
Figure 6 presents a summary of error types identified in the solution generation process of EHRAgent based on MIMIC-III dataset, as determined through manual examinations and analysis. The majority of errors occur because the LLM agent consistently fails to identify the root cause of these errors within $T$-step trails, resulting in plans that are either incomplete or inexcusable. Additional analysis of each error type is available in Appendix D.2.
@@ 5 Related Work
Augmenting LLMs with External Tools. LLMs have rapidly evolved from text generators into core computational engines of autonomous agents, with advanced planning and tool-use capabilities (Schick et al., 2023; Shen et al., 2023; Wu et al., 2023a; Karpas et al., 2022). LLM agents equip LLMs with planning capabilities (Wei et al., 2022; Yao et al., 2023a) to decompose a large and hard task into multiple smaller and simpler steps for efficiently navigating complex real-world scenarios. By integrating with external tools, LLM agents access external APIs for additional knowledge beyond training data (Shen et al., 2023; Huang et al., 2022; Sun et al., 2023; Lu et al., 2023; Patil et al., 2023; Qin et al., 2023b; Li et al., 2023), such as real-time information (Nakano et al., 2022; Parisi et al., 2022), computational capability (Schick et al., 2023), and coding proficiency (Wu et al., 2023b; Zhang et al., 2023; Gao et al., 2023; Chen et al., 2023a; Nan et al., 2023). The disconnection between plan generation and execution, however, prevents LLM agents from effectively and efficiently preventing error propagation and learning from environmental feedback (Yao et al., 2023b; Shinn et al., 2023; Yang et al., 2023). To this end, we leverage interactive coding to learn from dynamic interactions between the planner and executor, iteratively refining generated code by incorporating insights from error messages. Furthermore, EHRAgent extends beyond the limitation of short-term memory obtained from in-context learning, leveraging long-term memory (Sun et al., 2023; Wang et al., 2023b; Chen et al., 2023a; Zhuang et al., 2023a) by rapid retrieval of highly relevant and successful experiences accumulated over time.
LLM Agents for Scientific Discovery. Augmenting LLMs with domain-specific tools, LLM agents have demonstrated capabilities of autonomous design, planning, and execution in accelerating scientific discovery (Wang et al., 2023a; Xi et al., 2023; Zhao et al., 2023; Wang et al., 2023c; Cheung et al., 2023), including organic synthesis (Bran et al., 2023), material design (Boiko et al., 2023), and gene prioritization (Jin et al., 2023). In the medical field, MedAgents (Tang et al., 2023), a multi-agent collaboration framework, leverages role-playing LLM-based agents in a task-oriented multi-round discussion for multi-choice questions in medical entrance examinations. Similarly, Abbasian et al. (2023) develop a conversational agent to enhance LLMs using Langchain tools [^2] https://github.com/langchain-ai/langchain for general medical question and answering tasks. Different from existing LLM agents in medical and scientific domains, EHRAgent integrates LLMs with interactive code interface, targeting complex tabular tasks derived from real-world EHRs through autonomous code generation and execution.
@@ 6 Conclusion
In this study, we developed EHRAgent, an LLM agent equipped with an interactive code interface for complex tabular reasoning on real-world EHRs. By leveraging the emergent few-shot learning capabilities of LLMs, EHRAgent enables autonomous code generation and execution to address complicated clinical tasks, including database operations on EHRs with minimal demonstrations. Furthermore, we improve EHRAgent by interactive coding with execution feedback, along with a long-term memory mechanism, thereby effectively facilitating plan optimization for multi-step problem-solving. Our experiments on real-world EHR datasets demonstrate the advantages of EHRAgent over baseline LLM agents in autonomous coding and improved medical reasoning.
Limitation
EHRAgent holds considerable potential for positive social impact in a wide range of clinical tasks and applications, including but not limited to patient cohort definition, clinical trial recruitment, case review selection, and treatment decision-making support. One potential limitation is that while the framework of our proposed EHRAgent is broadly applicable to various scenarios, it currently relies on code generation for tool usage and problem-solving. Furthermore, the adaptation and generalization of EHRAgent in low-resource languages is constrained by the availability of relevant resources and training data. Additionally, given the demands for privacy, safety, and ethical considerations in real-world clinical settings, our goal is to further advance EHRAgent by mitigating biases and addressing ethical implications, thereby contributing to the development of responsible artificial intelligence for healthcare and medicine.

### v3

@@ 3.4 Rubber Duck Debugging via Error Tracing
The interaction ends either when a ‘TERMINATE’ signal appears in the generated messages or when $t$ reaches a pre-defined threshold of steps $T$.
@@ 4 Experiments
Dataset ($\rightarrow$) MIMIC-III eICU TREQS
Complexity Level ($\rightarrow$) I II III IV All I II III All I II III All
Methods ($\downarrow$) /Metrics ($\rightarrow$) SR. SR. CR. SR. SR. CR. SR. SR. CR.
w/o Code Interface
CoT (Wei et al., 2022) 29.33 12.88 3.08 2.11 9.58 38.23 26.73 33.00 8.33 27.34 65.65 11.22 9.15 0.00 9.84 54.02
Self-Consistency Wang et al. (2023d) 33.33 16.56 4.62 1.05 10.17 40.34 27.11 34.67 6.25 31.72 70.69 12.60 11.16 0.00 11.45 57.83
Chameleon (Lu et al., 2023) 38.67 14.11 4.62 4.21 12.77 42.76 31.09 34.68 16.67 35.06 83.41 13.58 12.72 4.55 12.25 60.34
ReAct (Yao et al., 2023b) 34.67 12.27 3.85 2.11 10.38 25.92 27.82 34.24 15.38 33.33 73.68 33.86 26.12 9.09 29.22 78.31
Reflexion Shinn et al. (2023) 41.05 19.31 12.57 11.96 19.48 57.07 38.08 33.33 15.38 36.72 80.00 35.04 29.91 9.09 31.53 80.02
w/ Code Interface
LLM2SQL (Nan et al., 2023) 23.68 10.64 6.98 4.83 13.10 44.83 20.48 25.13 12.50 23.28 51.72 39.61 36.43 12.73 37.89 79.22
DIN-SQL (Pourreza and Rafiei, 2023) 49.51 44.22 36.25 21.85 38.45 81.72 23.49 26.13 12.50 25.00 55.00 41.34 36.38 12.73 38.05 82.73
Self-Debugging Chen et al. (2024) 50.00 46.93 30.12 27.61 39.05 71.24 32.53 21.86 25.00 30.52 66.90 43.54 36.65 18.18 40.10 84.44
AutoGen (Wu et al., 2023) 36.00 28.13 15.33 11.11 22.49 61.47 42.77 40.70 18.75 40.69 86.21 46.65 19.42 0.00 33.13 85.38
EHRAgent (Ours) 71.58 66.34 49.70 49.14 58.97 85.86 54.82 53.52 25.00 53.10 91.72 78.94 61.16 27.27 69.70 88.02
Table 1: Main results of success rate (i.e., SR.) and completion rate (i.e., CR.) on MIMIC-III, eICU, and TREQS datasets. The complexity of questions increases from Level I (the simplest) to Level IV (the most difficult).
@@ 4.1 Experiment Setup
Tasks and Datasets. We evaluate EHRAgent on three publicly available structured EHR datasets, MIMIC-III (Johnson et al., 2016), eICU (Pollard et al., 2018), and TREQS (Wang et al., 2020) for multi-hop question and answering on EHRs. These questions originate from real-world clinical needs and cover a wide range of tabular queries commonly posed within EHRs. Our final dataset includes an average of 10.7 tables and 718.7 examples per dataset, with an average of 1.91 tables required to answer each question. We include additional dataset details in Appendix A.
@@
Tool Sets. To enable LLMs in complex operations such as calculations and information retrieval, we integrate external tools in EHRAgent during the interaction with EHRs. Our toolkit can be easily expanded with natural language tool function definitions in a plug-and-play manner. Toolset details are available in Appendix B.
Baselines. We compare EHRAgent with nine LLM-based planning, tool use, and coding methods, including five baselines with natural language interfaces and four with coding interfaces. For a fair comparison, all baselines, including EHRAgent, utilize the same (a) EHR metadata, (b) tool definitions, and (c) initial few-shot demonstrations in the prompts by default. We summarize their implementations in Appendix C.
@@
Evaluation Protocol. Following Yao et al. (2023b); Sun et al. (2023); Shinn et al. (2023), our primary evaluation metric is success rate, quantifying the percentage of queries the model handles successfully. Following Xu et al. (2023); Kirk et al. (2024), we further assess completion rate, which represents the percentage of queries that the model can generate executable plans (even not yield correct results). We categorize input queries into complexity levels (I-IV) based on the number of tables involved in solution generation. We include more details in Appendix A.2.
@@
Implementation Details. We employ GPT-4 (OpenAI, 2023) (version gpt-4-0613) as the base LLM model for all experiments. We set the temperature to 0 when making API calls to GPT-4 to eliminate randomness and set the pre-defined threshold of steps ($T$) to 10. Due to the maximum length limitations of input context in baselines (e.g., ReAct and Chameleon), we use the same initial four-shot demonstrations ($K=4$) for all baselines and EHRAgent to ensure a fair comparison. Appendix E provides additional implementation details with prompt templates.
@@ 4.2 Main Results
Table 1 summarizes the experimental results of EHRAgent and baselines on multi-tabular reasoning within EHRs. From the results, we have the following observations:
(1) EHRAgent significantly outperforms all the baselines on all three datasets with a performance gain of 19.92%, 12.41%, and 29.60%, respectively. This indicates the efficacy of our key designs, namely interactive coding with environment feedback and domain knowledge injection, as they gradually refine the generated code and provide sufficient background information during the planning process. Experimental results with additional base LLMs are available in Appendix F.1.
(2) CoT, Self-Consistency, and Chameleon all neglect environmental feedback and cannot adaptively refine their planning processes. Such deficiencies hinder their performance in EHR question-answering scenarios, as the success rates for these methods on three datasets are all below 40%.
(3) ReAct and Reflexion both consider environment feedback but are restricted to tool-generated error messages. Thus, they potentially overlook the overall planning process. Moreover, they both lack a code interface, which prevents them from efficient action planning, and results in lengthy context execution and lower completion rates.
(4) LLM2SQL and DIN-SQL leverage LLM to directly generate SQL queries for EHR question-answering tasks. However, the gain is rather limited, as the LLM still struggles to generate high-quality SQL codes for execution. Besides, the absence of the debugging module further impedes its overall performance on this challenging task.
(5) Self-Debugging and AutoGen present a notable performance gain over other baselines, as they leverage code interfaces and consider the errors from the coding environment, leading to a large improvement in the completion rate. However, as they fail to model medical knowledge or identify underlying causes from error patterns, their success rates are still sub-optimal.
4.3 Ablation Studies
Complexity Level ($\rightarrow$) I II III IV All
Methods ($\downarrow$) /Metrics ($\rightarrow$) SR. SR. CR.
EHRAgent 71.58 66.34 49.70 49.14 58.97 85.86
w/o medical information 68.42 33.33 29.63 20.00 33.66 69.22
w/o long-term memory 65.96 54.46 37.13 42.74 51.73 83.42
w/o interactive coding 45.33 23.90 20.97 13.33 24.55 62.14
w/o rubber duck debugging 55.00 38.46 41.67 35.71 42.86 77.19
Our ablation studies on MIMIC-III (Table 2) demonstrate the effectiveness of all four components in EHRAgent. Interactive coding [^3] For EHRAgent w/o interactive coding, we deteriorate from generating code-based to natural language-based plans and enable debugging based on error messages from tool execution. is the most significant contributor across all complexity levels, which highlights the importance of code generation in planning and environmental interaction for refinement. In addition, more challenging tasks benefits more from knowledge integration, indicating that comprehensive understanding of EHRs facilitates the complex multi-tabular reasoning in effective schema linking and reference (e.g., tables, columns, and condition values) identification. Detailed analysis with additional settings and results is available in Appendix F.2.
@@ 4.4 Quantitative Analysis
Effect of Question Complexity. We take a closer look at the model performance by considering multi-dimensional measurements of question complexity, exhibited in Figure 4. Although the performances of both EHRAgent and the baselines generally decrease with an increase in task complexity (either quantified as more elements in queries or more columns in solutions), EHRAgent consistently outperforms all the baselines at various levels of difficulty. Appendix G.1 includes additional analysis on the effect of various question complexities.
@@
Sample Efficiency. Figure 5 illustrates the model performance w.r.t.  number of demonstrations for EHRAgent and the two strongest baselines, AutoGen and Self-Debugging. Compared to supervised learning like text-to-SQL (Wang et al., 2020; Raghavan et al., 2021; Lee et al., 2022) that requires extensive training on over 10K samples with detailed annotations (e.g., manually generated corresponding code for each query), LLM agents enable complex tabular reasoning using a few demonstrations only. One interesting finding is that as the number of examples increases, both the success and completion rate of AutoGen tend to decrease, mainly due to the context limitation of LLMs. Notably, the performance of EHRAgent remains stable with more demonstrations, which may benefit from its integration of a ‘rubber duck’ debugging module and the adaptive mechanism for selecting the most relevant demonstrations.
4.5 Error Analysis
Figure 6 presents a summary of error types identified in the solution generation process of EHRAgent based on the MIMIC-III, as determined through manual examinations and analysis. The majority of errors occur because the LLM agent consistently fails to identify the underlying cause of these errors within $T$-step trails, resulting in plans that are either incomplete or inexcusable. Additional analysis of each error type is available in Appendix G.2.
@@ 4.6 Case Study
Figure 7 presents a case study of EHRAgent in interactive coding with environment feedback. The initial solution from LLM is unsatisfactory with multiple errors. Fortunately, EHRAgent is capable of identifying the underlying causes of errors by analyzing error messages and resolves multiple errors one by one through iterations. We have additional case studies in Appendix H.
Figure 7: Case study of EHRAgent harnessing LLMs in a multi-turn conversation with a code executor, debugging with execution errors through iterations.
@@ 5 Related Work
Augmenting LLMs with External Tools. LLMs have rapidly evolved from text generators into core computational engines of autonomous agents, with advanced planning and tool-use capabilities (Schick et al., 2023; Shen et al., 2023; Wang et al., 2024b; Yuan et al., 2024a, b; Zhuang et al., 2023). LLM agents equip LLMs with planning capabilities (Yao et al., 2023a; Gong et al., 2023) to decompose a large and hard task into multiple smaller and simpler steps for efficiently navigating complex real-world scenarios. By integrating with external tools, LLM agents access external APIs for additional knowledge beyond training data (Lu et al., 2023; Patil et al., 2023; Qin et al., 2024; Li et al., 2023b, a). The disconnection between plan generation and execution, however, prevents LLM agents from effectively and efficiently mitigating error propagation and learning from environmental feedback (Qiao et al., 2023; Shinn et al., 2023; Yang et al., 2023). To this end, we leverage interactive coding to learn from dynamic interactions between the planner and executor, iteratively refining generated code by incorporating insights from error messages. Furthermore, EHRAgent extends beyond the limitation of short-term memory obtained from in-context learning, leveraging long-term memory (Sun et al., 2023; Zhang et al., 2023) by rapid retrieval of highly relevant and successful experiences accumulated over time.
LLM Agents for Scientific Discovery. Augmenting LLMs with domain-specific tools, LLM agents have demonstrated capabilities of autonomous design, planning, and execution in accelerating scientific discovery (Wang et al., 2023a, c, 2024a; Xi et al., 2023; Zhao et al., 2023; Cheung et al., 2024; Gao et al., 2024), including organic synthesis (Bran et al., 2023), material design (Boiko et al., 2023), and gene prioritization (Jin et al., 2024). In the medical field, MedAgents (Tang et al., 2023), a multi-agent collaboration framework, leverages role-playing LLM-based agents in a task-oriented multi-round discussion for multi-choice questions in medical entrance examinations. Similarly, Abbasian et al. (2023) develop a conversational agent to enhance LLMs using external tools for general medical question-answering tasks. Different from existing LLM agents in the medical domains that focus on improving tasks like multiple-choice question-answering, EHRAgent integrates LLMs with an interactive code interface, exploring complex few-shot tabular reasoning tasks derived from real-world EHRs through autonomous code generation and execution.
@@ 6 Conclusion
In this study, we develop EHRAgent, an LLM agent with external tools for few-shot multi-tabular reasoning on real-world EHRs. Empowered by the emergent few-shot learning capabilities of LLMs, EHRAgent leverages autonomous code generation and execution for direct communication between clinicians and EHR systems. We also improve EHRAgent by interactive coding with execution feedback, along with accumulative medical knowledge, thereby effectively facilitating plan optimization for multi-step problem-solving. Our experiments demonstrate the advantages of EHRAgent over baseline LLM agents in autonomous coding and improved medical reasoning.
Limitation and Future Work
EHRAgent holds considerable potential for positive social impact in a wide range of clinical tasks and applications, including but not limited to patient cohort definition, clinical trial recruitment, case review selection, and treatment decision-making support. Despite the significant improvement in model performance, we have identified several potential limitations of EHRAgent as follows:
Additional Execution Calls.
We acknowledge that when compared to open-loop systems such as CoT, Self-Consistency, Chameleon, and LLM2SQL, which generate a complete problem-solving plan at the beginning without any adaptation during execution; EHRAgent, as well as other baselines that rely on environmental feedback like ReAct, Reflexion, Self-Debugging, and AutoGen, require additional LLM calls due to the multi-round conversation. However, such open-loop systems all overlook environmental feedback and cannot adaptively refine their planning processes. These shortcomings largely hinder their performance for the challenging EHR question-answering task, as the success rates for these methods on all three EHR datasets are all below 40%. We can clearly observe the trade-off between performance and execution times. Although environmental feedback enhances performance, future work will focus on cost-effective improvements to balance performance and cost (Zhang et al., 2023).
Translational Clinical Research Considerations.
Given the demands for privacy, safety, and ethical considerations in real-world clinical research and practice settings, our goal is to further advance EHRAgent by mitigating biases and addressing ethical implications, thereby contributing to the development of responsible artificial intelligence for healthcare and medicine. Furthermore, the adaptation and generalization of EHRAgent in low-resource languages is constrained by the availability of relevant resources and training data. Due to limited access to LLMs’ API services and constraints related to budget and computation resources, our current experiments are restricted to utilizing the Microsoft Azure OpenAI API service with the gpt-3.5-turbo (0613) and gpt-4 (0613) models. As part of our important future directions, we plan to enhance EHRAgent by incorporating fine-tuned white-box LLMs, such as LLaMA-2 (Touvron et al., 2023).
Completion Rate under Clinical Scenarios.
Besides success rate (SR) as our main evaluation metric, we follow Xu et al. (2023); Kirk et al. (2024) and employ completion rate (CR) to denote the percentage of queries for which the model can generate executable plans, irrespective of whether the results are accurate. However, it is important to note that a higher CR may not necessarily imply a superior outcome, especially in clinical settings. In such cases, it is generally preferable to acknowledge failure rather than generate an incorrect answer, as this could lead to an inaccurate diagnosis. We will explore stricter evaluation metrics to assess the cases of misinformation that could pose a risk within clinical settings in our future work.
Privacy and Ethical Statement
In compliance with the PhysioNet Credentialed Health Data Use Agreement 1.5.0444https://physionet.org/about/licenses/physionet-credentialed-health-data-license-150/, we strictly prohibit the transfer of confidential patient data (MIMIC-III and eICU) to third parties, including through online services like APIs. To ensure responsible usage of Azure OpenAI Service based on the guideline [^5] https://physionet.org/news/post/gpt-responsible-use, we have opted out of the human review process by requesting the Azure OpenAI Additional Use Case Form [^6] https://aka.ms/oai/additionalusecase, which prevents third-parties (e.g., Microsoft) from accessing and processing sensitive patient information for any purpose. We continuously and carefully monitor our compliance with these guidelines and the relevant privacy laws to uphold the ethical use of data in our research and operations.